A city population has been increasing 3% each year. If the initial population was 15,500, write the exponential function then determine how many people there are after 12 years?


1
Expert's answer
2021-03-10T08:58:00-0500

Solution:

y =C(1+r)x

Where C is Initial Population

r is growth rate

x is time interval


Given that initial population is 15,500

x = 0


So 15,500 = C(1+r)^0

15,500 =C

or C= 15,500 ----------------------- (1)


y = C(1+r)x

y= 15,500 (1+.03)12 -----------------(2)

y = 15,500(1.03)12

y = 15,500 x 1.4257

y = 22,098